Amortization schedule

a. Set up an amortization schedule for a $25,000 loan tobe repaid in equal installments

at the end of each of the next 3 years. The interest rateis 10 percent, compounded

annually.

b. What percentage of the payment represents interest andwhat percentage represents

principal for each of the 3 years? Why do thesepercentages change over time?
Answer #1

643-2-2.34P SA: 1810472

a. Given that I = 10%, PV = 25,000 and N = 3

PMT [1 – (1+i)-n/i] = 25,000

PMT [1 – (1+0.1)-3/0.1] = 25,000

PMT*2.4869 = 25,000

PMT = 25,000/2.4869

PMT= 10052.87

Year

Loan

Interest

Installment

Capital Repay

Capital outstanding at the end

1

25000

2500

10052.87

7552.87

17447.13

2

17447.13

1744.713

10052.87

8308.157

9138.973

3

9138.973

913.8973

10052.87

9138.973

0

b. Percentage of Interest Pay

1st = 2500 / 10052.87 x 100 = 24.87 %

2nd = 1744.713 /10052.87 x 100 = 17.36 %

3rd = 913.8973 / 10052.87 x 100 = 9.09 %

Percentage of Capital payment

1st = 7552.87 / 10052.87 x 100 = 75.13 %

2nd = 8308.157 / 10052.87 x 100 = 82.64 %

3rd = 9138.9727 / 10052.87 x 100 = 90.91 %

These percentages changes with time as interest rate declines.